OpenStudy (anonymous):

Graph the hyperbola with equation quantity x plus 2 squared divided by 49 minus the quantity of y plus 3 squared divided by 4 = 1.

OpenStudy (anonymous):

\[\frac{(x+2)^2}{49}-\frac{(y+2)^2}{4}=1\] ?